# Tailoring Chatbots for Higher Education: Some Insights and Experiences

### v1

@@ Abstract
The general availability of powerful Large Language Models had a powerful impact on higher education, yet general models may not always be useful for the associated specialized tasks. When using these models, oftentimes the need for particular domain knowledge becomes quickly apparent, and the desire for customized bots arises. Customization holds the promise of leading to more accurate and contextually relevant responses, enhancing the educational experience. The purpose of this short technical experience report is to describe what “customizing” Large Language Models means in practical terms for higher education institutions. This report thus relates insights and experiences from one particular technical university in Switzerland, ETH Zurich.
@@ I Large Language Models
Large Language Models (LLMs) are only one particular class of artificial intelligence systems, yet they have garnered enormous attention due to their ease-of-use, public availability, reasoning capabilities, and general, broad knowledge base. This impact was also felt in higher education [1, 2, 3]. Due to the in-depth and specialized nature of higher education, LLMs could frequently benefit from more in-depth knowledge of particular domains. The most frequently used way of accessing these models is through chatbot interfaces, and thus the term “custom chatbot” gained traction for specialized LLMs.
LLMs are sophisticated AI systems driven by numerous parameters, particularly weights, which play a crucial role in their operation. At the heart of LLMs is an auto-complete-like mechanism based on a neural network, which, based on a sequence of existing text and those weights predicts what should come next in the sequence. This is somewhat similar to the next-word suggestions on a smartphone when sending a text message, only that the algorithm operates on smaller units called tokens. These tokens, which can represent whole words, parts of words, individual characters or punctuation, are the building blocks the model uses to generate text. The network learns probable sequences of tokens from existing documents, its training corpus.
Generating novel text by an LLM is called inference. Typically, chatbots generate these token sequences in response to user input, so-called prompts. Due to the probabilistic nature of LLMs, these responses were initially derided by some skeptics (including the author) as nothing more than “plausible fiction,” but in the meantime, the systems have proven to be useful across a wide spectrum of applications. Inference involves large-scale matrix operations that are usually performed on processing units that were originally designed for graphics operations, GPUs. In addition to well-known chatbot interfaces, and of possibly equal importance to higher education institutions, are application programming interfaces (APIs), so other systems can make use of the LLM capabilities.
@@ II Customized Large Language Models
To customize Large Language Models, there are basically three routes with very different complexities.
@@ II.1 Training from Scratch
By far the most complex and resource-intensive way to get a customized LLM is to build a model from scratch. Besides having to decide on crucial architectural features, today’s models have billions of weight parameters, which require an enormous amount of training materials to properly adjust and optimize; the training corpuses of today’s general-purpose models encompass trillions of tokens.
II.1.1 Commercial Systems
@@
Common lore is that powerful commercial systems like GPT-4 have essentially ingested the “whole internet;” while that may be a hyperbole, curating the training corpus is extremely work-intensive, followed by computationally intensive training over months, followed by supervised and unsupervised tuning and detoxing. The origin of the training materials is sometimes questionable and the legal situation murky [4], the computing requires enormous resources and power (associated with large CO2 production [5]), and the manual tuning at times employs questionable work practices [6].
Due to the high associated cost (said to be in the billions of dollars for powerful foundation models like OpenAI/Microsoft’s GPT-4 [7], Google’s Gemini [8], Meta’s (maker of Facebook) Llama [9], or Apple Intelligence [10]), this remains largely the domain of large corporations, even though models made by smaller companies such as Mistral [11] and Anthropic’s Claude [12] have also shown remarkably high promise.
@@
Models like GPT-4o [13] and Gemini are also multimodal, so they accept auditory and visual input, and they can produce corresponding output. In the area of higher education, multimodality is particularly interesting for lecture transcriptions (voice and blackboard), as well as handwriting recognition for grading tasks [14, 15]. This generalized multimodality currently is the domain of the largest commercial models; university-generated systems usually implement multimodality only in research settings.
In short, training a competitive LLM from scratch seems inadvisable to individual higher education institutions; it appears virtually impossible for academic projects to reach the same level of general knowledge, reasoning, and multimodality that commercial models achieve.
II.1.2 Public Systems
Smaller, domain-specific LLMs can be generated on the scale of countries or university consortia, such as the efforts currently undertaken by our Swiss AI project [16]. Such a project requires a large repository of carefully curated training materials, which can be challenging to obtain while respecting author rights; the custom material alone would be insufficient to establish language and reasoning skills in the models.
For the education model in particular, there are publicly available datasets for “basic training,” that is, content for typical high school and introductory college curricula, but it turned out to be an uphill battle to obtain specialized and institution-specific course materials from faculty. There is a notable availability gap between basic content and research content, which is again readily available, for example from preprint servers or open-access publications.
On the flip side, a significant advantage of this approach is the control over training materials, which enhances the trustworthiness, transparency, and compliance of the resulting AI systems. The Swiss AI Initiative aims to develop advanced, large-scale AI models aligned with Swiss values, leveraging the country’s supercomputing capabilities to foster innovation and accessibility in AI technology. Throughout all its efforts, the initiative adheres to guiding principles of using region-specific datasets, respecting copyright, and honoring website restrictions regarding content processing. This approach not only ensures legal and ethical compliance but also contributes to the creation of AI systems tailored to Switzerland’s needs and values. Also here, training a model from scratch requires considerable super-computing resources; for example, Swiss AI will use a machine with 10,000 high-end NVIDIA GPUs, which recently came online and has yet to be used for this purpose. Once again, an ambitious effort, which should not be lightly undertaken.
This approach seems appropriate for specialized domains when large-scale resources are available. Since the resulting model will be non-mainstream, it needs to be clarified early on where the model should eventually run (inference).
@@ II.2 Fine-Tuning
Fine-tuning takes an already pre-trained model and modifies the weights of the neural net (in some approaches, this is only done with a subset of tokens in the later layers of the network). This is similar to editing a pre-written paper: the core content is already in place, and now it is being refined and enhanced. A big advantage is that the model already “knows” how to talk and reason, and one only needs to provide the custom data, alas usually after extensive preparation in specialized formats. This requires much less, but still considerable computational effort. After downloading the weights for the pre-trained model, fine-tuning usually requires several rounds of fine-tuning; there are common libraries for such efforts, and typical projects make use of the Hugging Face ecosystem [17].
With the notable exception of Llama 3, large, full-featured commercial models are not available as “open-weight” pre-trained models to start from for these independent projects. For Gemini and Mistral, however, smaller versions are available open-weight, and for Claude, older versions are open-weight. None of these smaller models have multimodal capabilities, and smaller models are usually worse for non-English languages [18] (this is an issue in Switzerland, which has four official non-English languages, one of them lesser spoken). Llama 3.1, on the other hand, is a full-featured open-weight model, available up to 405 billion parameters; it also has the capability of producing multimodal output.
While GPT allows for limited fine-tuning, this has to be done on OpenAI’s platform at considerable cost per token [19], and the resulting model needs to be run there; it also can be run there, which might be advantageous if no other inference platform is available.
Within the Swiss AI initiative, an educational model is currently in development, with a first version based on Llama 3.1. A key focus is on domain-specific fine-tuning, allowing for targeted performance improvements in particular areas of importance to Swiss society and industry.
@@
While the effort is reasonable, there are caveats. The open-weight models are already carefully pre-trained and tuned, and fine-tuning them can lead to overall worse results: the system may gain factual knowledge, but lose language and reasoning capabilities, and it will likely “forget” other things [20]. Finding the right level of fine-tuning is a balancing act [21]. It also should be kept in mind that fine-tuned models still “hallucinate,” which is the term commonly used for producing incorrect responses not founded in training materials. Also, to take advantage of newer versions of the open-weight model, for example going from Llama 3.0 to Llama 3.1, the fine-tuning would have to be repeated, as all of the weight adjustments would need to be recalculated.
@@
Fine-tuning, however, can help a model better “speak the language” of particular disciplines. The most viable route here seems to be using Llama and perform a limited amount of fine-tuning with carefully curated and prepared custom data on university research-computing infrastructure. The Swiss AI initiative will also follow this approach for some models. As discussed later, just like for models trained from scratch, this effort should not be undertaken without having arrangements for eventually running the model (inference).
@@ II.3 Augmentation
Some commercial models allow for straightforward but limited customizations, such as GPTs [22], where a limited number of documents, roles, and prompt-wrappers can be supplied. On the downside, this service requires users to have a subscription, and the amount of data is limited. Also, the documents would reside at OpenAI, which at least at our university limits them to those classified as public.
Very similar to this, Retrieval Augmented Generation (RAG) [23] is a method of customizing a chatbot without changing the LLM. The basic concept is to send relevant reference material alongside the user input: the user submits a prompt to the bot, the bot finds relevant text segments in its database of custom documents, and it then prompts the LLM along the lines of “Reply to [user prompt] using the following background materials: [relevant text segments].“ Additional items injected into the prompt or the role may contain instructions of how to deal with situations where no relevant information could be found.
@@
This approach entails creating a local infrastructure for RAG either through a few hundred lines of code [25, 26] on a local low-power server using standard tools like LangChain [27] (the local machine does not do any heavy computing) or buying this service from a cloud provider (“RAG as a service”). For local installations, the service also requires access to a standard LLM via its API as the conversation and reasoning agent [28]. Figure 1 gives an example of a dialogue with such a system.
@@
At higher education institutions, it makes sense to set up one bot instance per course, each one using its own database file or database collection within the shared RAG infrastructure. The one-bot-per-course approach ensures that answers are specific to that course and follow the established notations. Most of all, though, learner questions about Calculus 1 should not be answered using concepts from Calculus 3. Documents for each course (scripts, exercise sheets, syllabi, etc.) can be collected in a folder (instructors frequently deliver them in a ZIP file); at ETH Zurich, these are usually a few hundred pages across ten to twenty documents.
For the semantic search, these documents need to be embedded, which can take a few seconds for single documents to minutes for course scripts and hours for extensive databases (this includes artificial wait times, since API-access to the embeddings is usually subject to token-rate restrictions; we found it practical to submit ten chunks at a time with a two-second wait in-between). Once the embedding and/or indexing is finished, the bot instance can be started.
@@
RAG has many advantages, as it is fast to implement and using standard LLMs, so inference is readily available. The system can be “rewired” (even on-the-fly) to take advantage of different LLMs or switch to a newer version of the LLM; at ETH, we are considering building a “model garden.” It can also be used in combination with customized LLMs. If prompted to say “I don’t know” in case the documents provide no relevant information, it is low on hallucinations (but also potentially less useful).
III Inference
A frequently overlooked issue when discussing customized bots is: the associated LLM needs to run somewhere. For mainstream commercial models, inference services are readily available in the cloud. While several services are available as freebies after registration (user access to everyday model versions, developer access to APIs), the more advanced systems usually come with a cost: individual subscriptions or site-licenses [29] for advanced chatbot interfaces, usage-based or contingent-based payments-per-token for production APIs. In practice, for our RAG-based bot, it turned out that inference costs $7.50 per student per course per semester for Azure AI Services.
As a cloud service, privacy and data security concerns are an issue. These can be addressed by constructing a proxy that anonymizes the requests of chatbots (e.g. ProxyGPT [30] or HAWKI [31]) or using existing Azure contracts [32]. The latter also guarantees that interactions are not used for training, which is crucial, as intentionally or unintentionally confidential information is submitted to the bot by users.
Running inference for open-weight models also comes at a price. At our university, it is still an open question where to run inference for new models that have been trained from scratch or fine-tuned. While inference takes less compute-power than training, it still requires several GPUs in 24/7-operation. Most universities have access to supercomputing resources, but these systems are usually designed for batch operation, and it is a challenge to permanently take nodes out of research commission. Obtaining cloud-computing resources is an option, for example using AWS SageMaker [33], but this can be costly and has the same privacy concerns mentioned earlier.
IV Disclaimer
@@
This report was put together in the hopes that it may be useful. It can be no more than a snapshot in time, written from the perspective of a technical university on August 26, 2024. The LLM landscape, and in particular the landscape of associated services, is rapidly evolving: systems come and go at a pace that can make this report obsolete within months.
V Conclusion

### v2

@@ Abstract
The general availability of general-purpose Large Language Models continues to impact on higher education, yet they may not always be useful for specialized tasks. When using these models, oftentimes the need for particular domain knowledge becomes quickly apparent, and the desire for customized bots arises. Customization holds the promise of leading to more accurate and contextually relevant responses, enhancing the educational experience. This report relates insights and experiences from one particular technical university in Switzerland, ETH Zurich, to describe what “customizing” Large Language Models means in practical terms for higher education institutions.
@@ I Large Language Models
Large Language Models (LLMs) are just one class of artificial intelligence systems, yet they have garnered enormous attention across society, including higher education [1, 2, 3, 4, 5]. LLMs are driven by numerous parameters, typically billions of learned “weights.” Learning here is based on corpora of training materials, which are huge sets of human-produced documents gathered across repositories and the internet; what the system “knows” depends on the composition of these materials, where today’s general-purpose models aim for a balanced set of general knowledge, optimized for breadth over depth — a mile wide and an inch deep. In our university context, however, we often require in-depth knowledge within specific domains, which is why there are frequent calls for customized models among academics. This experience report from ETH Zurich aims to illustrate what customization at different levels entails.
At the heart of LLMs is an auto-complete-like mechanism that, given preceding text and these weights, predicts what should come next in the sequence. For illustration, consider a deliberately simplified, whole-word toy example with a tiny training corpus of two sentences:
The Alps are a large mountain range in Europe.
The Matterhorn is located in the Alps.
When given the prompt “One of the mountains in Europe …,” such a toy model might continue “…is the Matterhorn.” The final result is, “One of the mountains in Europe is the Matterhorn.” Note how it draws inferences between the two training sentences: the Matterhorn is in the Alps and the Alps are in Europe. The result is useful and true.
Now consider a different two-sentence training corpus:
The Alps are a large mountain range in Europe.
The Alps are ideal for skiing.
With that corpus, the model has seen the range, the Alps, but no specific mountain names. Given the same prompt, the toy model may somewhat clumsily produce “…is the Alps.” The final statement, “One of the mountains in Europe is the Alps,” sounds plausible but is incorrect: the Alps are a range, not a mountain. Lacking specific examples, the model falls back to the only salient entity it has seen. In general, the less specific and less relevant the training material, the more the model’s completions drift toward improbable or unsupported statements (so-called “hallucinations”). If the model were human, one might say it does not know what it is talking about.
Real LLMs operate on the level of so-called tokens, which are generally not whole words, but just parts of words, syllables, or punctuation. Our whole-word illustration above also omits important realities of modern LLMs (embeddings, long-range context, calibration, and instruction-following). The point here is narrower: the data the model was trained on strongly influences what it finds “natural” to say next, even for simple prompts. Different model families, such as GPT [6], Gemini [7], and LLaMA [8], differ in architecture, parameter count, and training corpora. As a rough rule of thumb, larger, more extensively trained models tend to perform better; however, advances in reasoning-oriented training, instruction tuning, and data curation have made size less determinative than it once was.
Performing these inference operations (“running the LLM”) involves large-scale matrix operations that are usually performed on processing units that were originally designed for graphics operations (GPUs). Given the size of today’s models, this is typically the domain of large-scale computing centers — it would generally not be economically feasible for a single university to run any of the more powerful models 24/7 in a scalable way that would for example sustain hundreds of students in a lecture hall hitting the system at the same time.
Chatbots such as OpenAI’s ChatGPT, which is an end-user interface for GPT models, are the most prominent application of LLMs. However, of possibly equal or higher importance to higher education institutions, are application programming interfaces (APIs), so other campus systems can make use of the LLM capabilities in the background. Those systems may well run on-campus and use LLM-inference as a commodity service.
@@ II Customized Large Language Models
To customize Large Language Models, there are basically three routes with very different complexities, illustrated in Fig. 1 and discussed below.
@@ II.1 Training from Scratch
By far the most complex and resource-intensive way to get a customized LLM is to build a model from scratch. Besides having to decide on crucial architectural features, today’s models have billions of weight parameters, which require an enormous amount of training data to properly adjust and optimize; the training corpora of today’s general-purpose models encompass trillions of tokens. Curating these training corpora is work-intensive, followed by computationally intensive training over months, followed by supervised and unsupervised tuning and and safety alignment (often called “detoxification”), which aims to reduce the likelihood of unsafe, unlawful, or harmful outputs and to mitigate the influence of problematic web content. The computing requires considerable resources and power (associated with large CO2 production [9]), The process is illustrated in the top row of Fig. 1.
II.1.1 Public Systems
ETH Zurich (ETHZ) and its sister institution, EPF Lausanne (EPFL), recently released Apertus [10], which is built from scratch by our SwissAI initiative [11]. Designed to be open and fair, it limited its training corpus to documents permitted for use. As already within its own borders, Switzerland has four official languages, the training corpus also had a higher-than-usual proportion of non-English documents added to the mix. This project was made possible by having access to the high-performance Alps computer [12] (not to be confused with the European mountain range that contains the Matterhorn!). Alps is the Swiss National Supercomputing Centre’s HPE Cray EX254n system built with NVIDIA GH200 Grace Hopper Superchips, delivering 434.9 PFLOPS for AI- and data-intensive science and ranked #8 on the June 2025 TOP500 list of most powerful computers — all this techno-babble to illustrate that training a model from scratch requires computing resources that are usually only available on a national level.
Such a high-performance computer is designed for research, not production purposes. It processes in batches instead of being available 24/7, and one cannot dedicate all its compute power to just one specific application. The day-to-day grind of running the LLM, that is, performing inference, has to be done elsewhere, and that is the next challenge: you need a system that continuously operates at high availability. Without inference infrastructure, training such a model is like building a car in a world that has no roads. You can run it for benchmarks on your own test track, but it never goes anywhere.
Loading and unloading a specific model takes longer than the average user has patience to wait for a reply, so you need to run it “hot,” which requires permanently dedicating sufficient GPU resources to absorb peak loads, This is usually the domain of commercial hyperscalers like Microsoft’s and Google’s data centers. To make their efforts worthwhile and commercially viable, they tend to dedicate their computer resources to the usual suspects of general-purpose, commercial models, and it is hard for a from-scratch, open model to break into this (and if the hyperscalers load such models, it is usually by-the-hour, not per usage: you pay for the time that they dedicate their hardware for your purposes, not for how much it is actually used).
However, besides building another Large Language Model, in specialized applications, Small Language Models might do the trick (or even perform better). As the name suggests, these models are much smaller, so they require less compute-power in training, and their inference requirements are far more modest — in exchange for their knowledge being an inch wide and a mile deep. Coupled with reasoning abilities, these are serious contenders to LLMs in their respective domains.
II.1.2 Commercial Systems
@@
Common lore is that powerful commercial systems like GPT-5 [13, 14] have essentially ingested the “whole internet;” while that may be a hyperbole, the origin of the training materials is sometimes questionable and the legal situation murky [15]. Also, the manual tuning and detoxification at times employs questionable work practices [16]. As a result, though, their performance is hard to achieve with more ethically built models, and in spite of concerns, many of us rely on their convenience and ever-increasing reliability.
@@
In addition, models like GPT-4o [17] , GPT-5, and Gemini are multimodal, so they accept auditory and visual input, and they can produce corresponding output. In the area of higher education, multimodality is particularly interesting for lecture transcriptions (voice and blackboard), as well as handwriting recognition for grading tasks [18, 19]. This generalized multimodality currently is the domain of the largest commercial models; university-generated systems usually implement multimodality only in research settings.
Inference for commercial models is readily available via managed APIs and is typically billed on a usage basis (with some enterprise/education plans also charging per seat; using the earlier analogy, there is plenty of road for these cars). For European universities, major vendors offer contractual terms that exclude customer prompts and completions from model training and provide EU data-residency options or EU processing paths. From an environmental standpoint, shared cloud capacity tends to keep GPUs highly utilized (which reduces power consumption per workload), and large providers increasingly site or procure power near low-carbon grids. For example, ETH Zurich currently uses Swedish data centers via Microsoft Azure contracts.
The trade-off is reduced transparency and control over training-data provenance, data governance, and labor practices in alignment work; institutions effectively purchase availability, cost efficiency, and convenience at the expense of some aspects of digital sovereignty. We will see later how these models can be customized.
@@ II.2 Fine-Tuning
Fine-tuning takes an already pre-trained model and modifies the weights. This only works if the original model is “open-weight,” meaning, those billions of numbers can be downloaded and locally modified (our Apertus is such an open-weight model; while GPT allows for limited fine-tuning (illustrated in the middle row of Fig. 1), this has to be done on OpenAI’s platform at considerable cost per token [20], and the resulting model needs to be run there). This is similar to editing a pre-written paper: the core content is already in place, and now it is being refined and enhanced. A big advantage is that the model already “knows” how to talk and reason, and one only needs to provide the custom data, alas usually after extensive preparation in specialized formats. This requires much less, but still considerable computational effort. After downloading the weights for the pre-trained model, fine-tuning usually requires several iterations; there are common libraries for such efforts, and typical projects make use of the Hugging Face ecosystem [21]. The process is illustrated in the second-to-the-top row of Fig. 1.
@@
While the effort is reasonable, there are caveats. The open-weight models are already carefully pre-trained and tuned, and fine-tuning them can lead to overall worse results: the system may gain factual knowledge, but lose language and reasoning capabilities, and it will likely “forget” other things [22]. In our initial, simplified example, you could add the information that the Matterhorn is in the Alps, but if you are not careful, it may loose the association between the Alps and Europe. Finding the right level of fine-tuning is a balancing act [23]. It also should be kept in mind that fine-tuned models still “hallucinate.” Also, to take advantage of newer versions of the open-weight model, for example going from LLaMa 3.1 to LLaMa 3.2, the fine-tuning would have to be repeated, as all of the weight adjustments would need to be recalculated.
@@
Fine-tuning, however, can help a model better “speak the language” of particular disciplines. The most viable route here seems to perform a limited amount of fine-tuning with carefully curated and prepared custom data on university research-computing infrastructure. For example, a research group at ETH Zurich fine-tuned a LLaMa to better speak the language of mathematics tutors [24].
The finished product, unfortunately, has the same inference challenge: where do you run it? Just like with models that are trained from scratch, production requires dedicated hardware, and the more boutique models are being created, the bigger the challenge. There are mechanisms to only load and unload the changes between different fine-tunings of the same pre-trained model, thus reducing wait times when switching models, but that still requires inference infrastructure. Unless the goal is purely research (which of course is legitimate!), any proposal to fine-tune LLMs should include a section where to run the finished product.
@@ II.3 Augmentation
Retrieval Augmented Generation (RAG) [25] is a method of customizing a chatbot without changing the LLM. The basic concept is to send relevant reference material alongside the user input: the user submits a prompt to the bot, the bot finds relevant text segments in its database of custom documents, and it then prompts the LLM along the lines of “Reply to [user prompt] using the following background materials: [relevant text segments].“ For example, if the sentence “the Matterhorn is located in the Alps” were part of the text segments sent alongside the user prompt, even the model that out-of-the-box only “knows” about the Alps as a skiing place would be able to provide the correct inference. Additional items injected into the prompt or the role may contain instructions of how to deal with situations where no relevant information could be found, and whether or not to have the references that are sent along supersede general knowledge of the model. The setup is illustrated in the second-to-bottom row of Fig. 1
@@
This approach entails creating a local infrastructure for RAG either through a few hundred lines of code [27] on a local low-power server using standard tools like LangChain [28] (the local machine does not do any heavy computing) or buying this service from a cloud provider (“RAG as a service”). For local installations, the service also requires access to a standard LLM via its API as the conversation and reasoning agent [29]. Figure 2 gives an example of a dialogue with such a system.
@@
At higher education institutions, it makes sense to set up one bot instance per course, each one using its own database file or database collection within the shared RAG infrastructure. The one-bot-per-course approach ensures that answers are specific to that course and follow the established notations. Most of all, though, learner questions about Calculus 1 should not be answered using concepts from Calculus 3. Instructors can upload and remove documents for their courses, such as scripts, slides, old exams, exercise sheets; at ETH Zurich, these usually amount to several hundred pages. For the semantic search, these documents need to be embedded, so it usually takes a few minutes until new materials are available in the bot.
@@
RAG has many advantages, as it is fast to implement and using standard LLMs, so inference is readily available. The system can be “rewired” (even on-the-fly) to take advantage of different LLMs or switch to a newer version of the LLM; it can also be used in combination with customized LLMs (illustrated by the orange dashed lines in Fig. 1). If prompted system-side to say “I don’t know” in case the documents provide no relevant information, it is low on hallucinations (but also potentially less useful). It can also be prompted system-side to not directly answer leaner questions, but for example ask follow-up questions and guide the learner toward a solution.
Finally, this architecture can be expanded, and the same materials used to augment other course-specific functionality, such as practice problem generation [30], homework feedback [29], or exam grading [31, 32, 33].
The bottom row of Fig. 1 shows the simplest customization method, using OpenAI GPTs [34]. The technology is proprietary, but likely RAG. As opposed to local installations, the amount of reference material you can upload is limited, and you have no control on which data center inference is run. A subscription is required to generate GPTs, and unsubscribed users only have limited access to the GPTs before it reverts to a very simple (by today’s standards) LLM in the background.
III Disclaimer
@@
This report was put together in the hopes that it may be useful. It can be no more than a snapshot in time, written from the perspective of a technical university on October 26, 2025. The LLM landscape, and in particular the landscape of associated services, is rapidly evolving: systems come and go at a pace that can make this report obsolete within months.
IV Conclusion